# MAPLE: Enhancing Review Generation with Multi-Aspect Prompt LEarning in Explainable Recommendation

### v1

@@ Abstract
Explainable Recommendation task is designed to receive a pair of user and item and output explanations to justify why an item is recommended to a user. Many models treat review-generation as a proxy of explainable recommendation. Although they are able to generate fluent and grammatical sentences, they suffer from generality and hallucination issues. We propose a personalized, aspect-controlled model called Multi-Aspect Prompt LEarner (MAPLE), in which it integrates aspect category as another input dimension to facilitate the memorization of fine-grained aspect terms. Experiments on two real-world review datasets in restaurant domain show that MAPLE outperforms the baseline review-generation models in terms of text and feature diversity while maintaining excellent coherence and factual relevance. We further treat MAPLE as a retriever component in the retriever-reader framework and employ a Large-Language Model (LLM) as the reader, showing that MAPLE’s explanation along with the LLM’s comprehension ability leads to enriched and personalized explanation as a result. We will release the code and data in this http upon acceptance.
@@
Ching-Wen Yang, Che-Wei Chen, Kun-Da Wu, Hao Xu, Jui-Feng Yao, Hung-Yu Kao Department of Computer Science and Information Engineering, National Cheng-Kung Univeristy Google {P76114511, Q56104076}@gs.ncku.edu.tw hykao@mail.ncku.edu.tw {harrisonwu, hxx, eagley}@google.com
@@ 1 Introduction
In the context of Natural Language Generation (NLG) explainable recommendation models, the characteristics of a good explanation are 1) diverse: for the same item, a model should generate personalized rationales for different users 2) factual: the recommended feature or content should be factually relevant to the item 3) precise: the recommended feature should be as precise (as opposed to general) as possible, etc. In these criteria, we witness a "diversity-factuality dilemma." While models generating generic aspects like "food" in restaurants may offer safe, broadly relevant explanations, they risk being uninformative and repetitive (Li et al., 2017; Dong et al., 2017; Li et al., 2021, 2023b). On the other hand, models creating diverse and creative aspects produce more engaging explanations but may compromise factual relevance by including irrelevant details. Few existing models effectively balance these traits. To address this, we introduce the concept of "multi-aspect" from multi-aspect sentiment analysis (Lu et al., 2011; Xianghua et al., 2013), which assumes a limited inventory of multiple, distinct aspects like “food quality” or “service” in restaurant domain, providing more fine-grained understanding of user preferences. In this paper, we propose a simple yet effective prompt tuning approach that integrates aspect as an auxiliary signal to improve the memorization of rich aspect terms, called MAPLE (Multi-Aspect Prompt LEarner). We further control its factual relevance by incorporating the retriever-reader framework as its downstream. Our contributions are as follows:
@@
    MAPLE increases the diversity and precision of generated features, endorsed by the self-crafted explainability metrics from aspect-wise perspectives.
@@
    We renew the review-generation datasets in restaurant domain (Li et al., 2020) to include higher quality aspect terms and additionally label associated aspect categories, enhancing the research in this field.
@@ 2.1 Review Generation
Past works (Dong et al., 2017; Li et al., 2017, 2021, 2023b) have proposed end-to-end frameworks for learning a short explanation (often part of the review) for a user-item pair. Experimental statistics and case studies show that these models generate repetitive, overly-generic sentences that are still far from good rationales (Xie et al., 2023), and even suffer from hallucination issue (Xie et al., 2023; Maynez et al., 2020)  [^1] A model exhibiting hallucinations generates content includes inaccuracies or irrelevant information.. It is also observed that the uninformative sentences are often due to models’ being incapable of generating precise and informative aspect terms.
@@ 2.2 Aspect-aware Explanation Generation
According to Zhang et al. (2023) which studies into aspect-based sentiment analysis, aspect categories $c$ refer to broad attributes (such as food or service in the restaurant domain) and aspect terms $a$ refer to specific targets (such as the beef tacos in the review "The beef tacos here are amazing"). To alleviate the issue of overly general aspect terms, Ni and McAuley (2018) proposes ExpansionNet that expands short phrases into detailed explanation using a tri-encoder framework, in which it relies on the aspect encoder to predict aspect-term importance. To get the aspect terms, ExpansionNet uses an unsupervised aspect extraction model called ABAE (He et al., 2017) for inferring unsupervised aspect categories and the top-K aspect terms.
Building on ExpansionNet, UARM (Sun et al., 2021) enhances the use of inferred aspect embeddings from ABAE by converting them into aspect-aware user and item representations. Li et al. (2020) extracts review aspect terms with Sentires (Zhang et al., 2014), and its proposed model NETE does aspect-term condition generation in training. In inference stage, NETE uses Point-wise Mutual Information (PMI) to rank and select an aspect term by user preference. However, aspect-term distribution might be too sparse to be captured by PMI. Therefore, there’s an emerging line of works that simply assume the ground-truth aspect terms are given: Li et al. (2021) introduces PETER+ that places aspect-term tokens behind user and item IDs to guide the generation of explanations. UCEpic (Li et al., 2023a) employs a bi-directional transformer (Vaswani et al., 2017) that starts from the aspect term and iteratively predict and insert mask tokens, formulating the explanation process as a progressive text generation.
Our approach aligns more to the line of NETE, where in training stage, MAPLE is trained by a given aspect-category signal, and in inference stage, it decides which aspect signal it wants to use. MAPLE is a simple yet effective approach for controlling from a broad sense and generating fine-grained aspect terms.
@@ 2.3 Retrieval-Augmented Explanation Generation
To address the issue of factual relevance, a long-existing challenge in NLG task, Xie et al. (2023) proposes a retriever-reader two-stage framework called PRAG. In the first stage, a personalized retriever formulates a latent query based on the input user and item; in the second stage, a reader model generates the explanation, grounding it in the retrieved content, which consists of past reviews that are inherently factual. Given the current advancements in LLM, we argue that LLMs fit into the role of reader very well as their reading comprehension ability can be leveraged not only to enhance factual relevance but also to refine the style of the generated reviews, making them more persuasive and closely aligned with real explanations. In our case studies in Appendix D, with the examples of employing MAPLE as a discrete retriever and LLM as a reader, we showcase the potential for improving both factual accuracy and the persuasiveness of explanations.
@@ 3.1 Problem Setup
Consider a set of users $U$ and items $I$. For each user $u\in U$ and item $i\in I$ in the training set, there is an associated review $r$ covering several aspects $c_{i}$, each provides an explanation in terms of that aspect. The task of joint aspect-recommendation-explanation is to learn a function $rec:(u,i)\rightarrow(\hat{c}_{u,i},\hat{E}_{u,i})$, where $\hat{c}_{u,i}$ is the predicted aspect-category distribution $\in\mathcal{R}^{n_{\text{aspect}}}$ ($n_{\text{aspect}}$ denotes the number of aspect categories) and $\hat{E}_{u,i}$ is the textual explanation given to justify the recommendation of item $i$ to user $u$. In this context, $\hat{c}_{u,i}$ serves as a by-product that is used in assisting the generation of $\hat{E_{u,i}}$ in inference stage.
Figure 2: The proposed MAPLE architecture. MAPLE is trained on Task 1: Explanation Generation and Task 2: Aspect Recommendation as a unified architecture. During inference, MAPLE exploits the trained Aspect-Recommendation Model (Rec. Model) in Task 2 to predict an aspect distribution $\hat{c}_{u,i}$ and obtain an aspect signal $\hat{c}$, and then feed it back to Task 1 for generation.
@@ 3.2 Multi-Aspect Review Segmentation
Instead of predicting aspect-term importance as done by ExpansionNet and UARM, we decide to just let our model adopt a broader perspective by learning aspect topics and extracting the aspect terms from within the learned aspect topic, based on the ID information. Our rationale is that, the frequency of updates for any specific user or item during training is minimal in recommendation datasets. This limitation prevents the model from adequately learning the characteristics of individual users or items, leading to overly general explanations conditioning on IDs alone. By introducing aspect prompts $c$, the model focus on learning the fine-grained aspect terms of the particular category, while the ID embeddings are learned as selectors for relevant aspects. Once $c$ is robustly represented, the model effectively recommends aspect terms aligning with user interests and item features, enhancing explanation specificity and diversity.
3.3 Automated Sentiment-Analysis Pipeline
We employ an automated sentiment-analysis pipeline to extract sentiment tuples (aspect term, opinion, sentiment, category) from reviews. Concretely, we use a fine-tuned T5 (Raffel et al., 2020) to extract the tuples within the reviews. Reviews without sentiment tuples are discarded. We then use another fine-tuned BART-based classifier (Lewis et al., 2020) to assign each extracted aspect term with an aspect category (for details on model and aspect inventory, see Appendix A.1 and A.3). Finally, for every user-item pair, we check if there are multiple mined aspect terms under the same aspect category; if yes, we choose the tuple with an arbitrary select strategy, e.g., choose the one that is the longest, and add the associated tuple into the training data. Figure 1 illustrates an example review and the extracted sentiment tuples, since we take one tuple for each category, this review yields two tuples. For "Gluten-Free, Vegan, Vegetarian" category, we take ("vegan pizza", …) since it has a longer aspect term compared to "pie"; for "Seafood" category, we take ("chowder", …).
3.4 MAPLE: Aspect-based Explanation Generation
@@
3.5 MAPLE: Aspect Recommendation
During training, MAPLE takes a user-item pair and the auxiliary aspect category, in total three prompt tokens as its input. However, in inference stage, it is possible that the aspect information is absent. To close the gap between training and inference, we design an auxiliary task called aspect recommendation, to recover the relationship between IDs and their associated aspect categories. The input of aspect-recommendation task is a pair of user, item IDs, the output is the predicted aspect category probabilities $\hat{c_{u,i}}\in\mathbb{R}^{n_{\text{aspect}}}$. We implement the model with Multi-layer Perceptron (MLP):
@@ 3.5 MAPLE: Aspect Recommendation
where $\mathbf{W}_{h}\in\mathbb{R}^{h\times 2d}$ and $\mathbf{b}_{h}\in\mathbb{R}^{h}$ are the weights and biases of fully-connected layer(s),   $\mathbf{W}_{c}\in\mathbb{R}^{n_{\text{aspect}}\times h}$ and $\mathbf{b}_{c}\in\mathbb{R}^{n_{\text{aspect}}}$ are the weights and biases of the output classifier layer, and $\delta$ and $\sigma$ are the activation functions ReLU and Sigmoid, respectively. However, the aspect-category label exhibits a very skewed, long-tail distribution (Figure 4) poses a challenge in training the aspect-recommendation model, since a conventional classifier easily overfits to the head classes. To address this issue, we employ a specialized loss called distribution-balanced loss function (Wu et al., 2020). This approach considers both class-level and instance-level sampling frequencies to mitigate the imbalance in the multi-label scenario; it also counteracts the limitations of the sigmoid function, which treats each class independently and can cause overfitting for head classes. We observe that training with distribution-balanced loss effectively captures aspect preferences based on ID information, rather than the label distribution. The distribution-balanced loss is formally described as:
$\Delta_{i}=z^{k}_{i}-\nu_{i}$ (3)
@@
where $\mathcal{L}_{DB}(x^{k},y^{k})$ represents the distribution-balanced loss function for the $k$-th sample; $\hat{r}^{k}_{i}$ represents the sampling weighting factor for the $i$-th aspect class of the $k$-th sample; $y^{k}_{i}$ is the ground truth label for the $i$-th class of the $k$-th sample; $z^{k}_{i}$ is the logit for the $i$-th class of the $k$-th sample; $\nu_{i}$ is the class bias for $i$-th class, and $\lambda$ is a hyperparameter regularizing sigmoid’s gradient issue.
3.6 Model Training and Inference
Our objective is to train both the aspect-based explanation generation and aspect recommendation components simultaneously in an end-to-end manner.
Two-stage Training Approach. To integrate the aspect recommendation task, we need ID embeddings to be representative so that it predicts an aspect distribution, therefore we propose a two-stage training approach. Initially, we focus on the explanation generation task alone, optimizing it using negative log-likelihood $\mathcal{L}_{T}$ until convergence. During this phase, the ID embeddings adapt by learning aspect information from the co-occurring aspect category labels and text data. Once the ID embeddings become stable and representative, we incorporate the recommendation loss with weight $\alpha$ into the total loss to endow the ID embeddings the selector ability (see Equation 5).
$L=L_{T}+\alpha L_{DB}$ (5)
This two-stage tuning approach ensures that ID embeddings are robust and informative before they are used to predict aspect distributions, and proves that using one set of ID embeddings can effectively multi-task in terms of high-quality explanation generation and aspect-recommendation. Note that in training, we do not use the predicted aspect distribution but simply include it for loss calculation.
Inference. Given user and item IDs, we first inference Task 2’s aspect-recommendation model and trim the predicted aspect distribution $\hat{c}_{u,i}$ to leave only the top 5, and then we sample $K(K\leq 5)$ aspects with replacement from the trimmed distribution, using the predicted probabilities as weights. The $K$ embeddings are then fused (by taking its mean) to form the aspect signal at position $\hat{c}$, as shown in Figure 2 (b). The pretrained-transformer is thus conditioned on $u,i,\hat{c}$ to generate a well-rounded explanation $\hat{E}_{u,i}$.

### v2

@@ Abstract
The Explainable Recommendation task is designed to receive a pair of user and item and output explanations to justify why an item is recommended to a user. Many models approach review generation as a proxy for explainable recommendations. While these models can produce fluent and grammatically correct sentences, they often lack precision and fail to provide personalized, informative recommendations. To address this issue, we propose a personalized, aspect-controlled model called Multi-Aspect Prompt LEarner (MAPLE), which integrates aspect category as another input dimension to facilitate memorizing fine-grained aspect terms. Experiments conducted on two real-world review datasets in the restaurant domain demonstrate that MAPLE significantly outperforms baseline review-generation models. MAPLE excels in both text and feature diversity, ensuring that the generated content covers a wide range of aspects. Additionally, MAPLE delivers good generation quality while maintaining strong coherence and factual relevance. The code and dataset used in this paper can be found here  [^1] https://github.com/Nana2929/MAPLE.git.
@@
Ching-Wen Yangthanks: P76114511@gs.ncku.edu.tw, Zhi-Quan Fengthanks: P78123011@gs.ncku.edu.tw, Ying-Jia Linthanks: yjlin@cgu.edu.tw, Che-Wei Chenthanks: Q56104076@gs.ncku.edu.tw, Kun-Da Wu thanks: harrisonwu@google.com, Hao Xuthanks: hao.xu@nyu.edu, Jui-Feng Yaothanks: eagley@google.com, Hung-Yu Kaothanks: hykao@cs.nthu.edu.tw Department of Computer Science and Information Engineering, National Cheng-Kung University, Google Pixel Software
@@ 1 Introduction
In the context of Natural Language Generation (NLG) explainable recommendation models, a good explanation is required to have the following characteristics: 1) Diversity: for the same item, a model should generate personalized rationales for different users. 2) Factuality: the recommended feature or content should be factually relevant to the item. 3) Precision: the recommended feature should be as precise (as opposed to general) as possible.
Figure 1: Overview of explainable recommendation: A simple illustration of the "diversity-factuality dilemma".
Within these criteria, we observe a "diversity-factuality dilemma", as illustrated in Figure 1. Models that generate generic aspects, such as "food" in the context of restaurants, may provide safe and broadly applicable explanations but risk being uninformative and repetitive, which has been noted in prior studies (Li et al., 2017; Dong et al., 2017; Li et al., 2021, 2023). On the other hand, models creating diverse and creative aspects produce more engaging explanations but may compromise factual relevance by including irrelevant details, which is referred to as the hallucination problem. Few existing models effectively balance these traits.
To address this, we introduce the concept of "multi-aspect" from multi-aspect sentiment analysis (Lu et al., 2011; Xianghua et al., 2013), which assumes a limited inventory of multiple, distinct aspects like “food quality” or “service” in the restaurant domain, providing a more fine-grained understanding of user preferences. In this paper, we propose a simple yet effective two-stage tuning approach that integrates aspect as an auxiliary signal to improve the memorization of rich aspect terms, called MAPLE (Multi-Aspect Prompt LEarner). For diversity, we utilize a Multilayer Perceptron (MLP) model to predict aspects and employ a distribution-balanced loss function (Wu et al., 2020) to ensure diverse predictions. For factuality, we generate recommendation text using a retriever-reader framework, and for precision, we retrieve specific information to guide the LLM in generating user- and item-specific recommendation text. Our contributions are as follows:
@@
    MAPLE increases diversity, factuality, and precision of generated features, endorsed by the self-crafted explainability metrics from aspect-wise perspectives.
@@
    We renew the review-generation datasets in the restaurant domain (Li et al., 2020) to include higher quality aspect terms and additionally label associated aspect categories, enhancing the research in this field.
We treat MAPLE as a retriever in the retriever-reader framework with an LLM as the reader in Appendix D, demonstrating that MAPLE’s explanations combined with the LLM’s comprehension yield enriched, personalized results.
@@ 2.1 Review Generation
Past works (Dong et al., 2017; Li et al., 2017, 2021, 2023) have proposed end-to-end frameworks for learning a short explanation (often part of the review) for a user-item pair. Experimental statistics and case studies show that these models generate repetitive, overly generic sentences that are still far from good rationales (Xie et al., 2023), and even suffer from hallucination issue (Xie et al., 2023; Maynez et al., 2020)  [^2] A model exhibiting hallucinations generates content that includes inaccuracies or irrelevant information.. It is also observed that the uninformative sentences are often due to models’ being incapable of generating precise and informative aspect terms.
@@ 2.2 Aspect-aware Explanation Generation
According to Zhang et al. (2023) which studies into aspect-based sentiment analysis, aspect categories $c$ refer to broad attributes (such as food or service in the restaurant domain) and aspect terms $a$ refer to specific targets (such as the beef tacos in the review "The beef tacos here are amazing").
To address overly general aspect terms, Ni and McAuley (2018) proposed ExpansionNet, which expands short phrases into detailed explanations via a tri-encoder framework that predicts aspect-term importance using an aspect encoder. Aspect terms are obtained through ABAE (He et al., 2017), an unsupervised model that infers aspect categories and retrieves top-K aspect terms.
Building on ExpansionNet, UARM (Sun et al., 2021) enhances the use of inferred aspect embeddings from ABAE by converting them into aspect-aware user and item representations. Li et al. (2020) extracts review aspect terms with Sentires (Zhang et al., 2014), and its proposed model NETE does aspect-term condition generation in training. In the inference stage, NETE uses Point-wise Mutual Information (PMI) to rank and select an aspect term by user preference. However, aspect-term distribution might be too sparse to be captured by PMI. Therefore, there’s an emerging line of works that simply assume the ground-truth aspect terms are given: One of them is PETER+ (Li et al., 2021), which places aspect-term tokens behind user and item IDs to guide the generation of explanations. Another example is ERRA (Cheng et al., 2023), which leverages Sentence-BERT (Reimers and Gurevych, 2019) to encode aspect terms and the corpus. It searches for appropriate aspect and corpus vectors to support the text generation.
Our approach aligns more with the line of NETE, wherein in the training stage, MAPLE is trained by a given aspect-category signal, and in the inference stage, it decides which aspect signal it wants to use. Compared to ERRA, MAPLE fine-tunes the generation model and incorporates multi-aspect learning to predict aspects with greater diversity. This makes MAPLE a simple yet effective approach for controlling from a broad sense and generating fine-grained aspect terms.
@@ 2.3 Retrieval-Augmented Explanation Generation
To address the issue of factual relevance, a long-existing challenge in NLG task, Xie et al. (2023) proposes a retriever-reader two-stage framework called PRAG. In the first stage, a personalized retriever formulates a latent query based on the input user and item; in the second stage, a reader model generates the explanation, grounding it in the retrieved content, which consists of past reviews that are inherently factual. Given the current advancements in LLM Rorseth et al. (2024), we argue that LLMs fit into the role of the reader very well as their reading comprehension ability can be leveraged not only to enhance factual relevance but also to refine the style of the generated reviews, making them more persuasive and closely aligned with real explanations. In our case studies in Appendix D, with the examples of employing MAPLE as a discrete retriever and LLM as a reader, we showcase the potential for improving both factual accuracy and the persuasiveness of explanations.
@@ 3 Methodology
In this work, we propose a two-stage training approach. We initially extract the aspect term from reviews in Section 3.2 to support the two-stage training. In Stage 1 (Section 3.3), we focus exclusively on the explanation generation task, optimizing it using the negative log-likelihood loss $\mathcal{L}_{T}$ until convergence. In Stage 2 (Section 3.4), we introduce the recommendation loss, weighted by $\alpha$, into the total loss to equip the ID embeddings with the selector ability. Finally, our inference process is presented in Section 3.5.
@@ 3.1 Problem Setup
Consider a set of users $U$ and items $I$. For each user $u\in U$ and item $i\in I$ in the training set, there is an associated review $r$ covering several aspects $c_{i}$, each explained in terms of that aspect. The task of joint aspect-recommendation-explanation is to learn a function $rec:(u,i)\rightarrow(\hat{c}_{u,i},\hat{E}_{u,i})$, where $\hat{c}_{u,i}$ is the predicted aspect-category distribution $\in\mathcal{R}^{n_{\text{aspect}}}$ ($n_{\text{aspect}}$ denotes the number of aspect categories) and $\hat{E}_{u,i}$ is the textual explanation given to justify the recommendation of item $i$ to user $u$. In this context, $\hat{c}_{u,i}$ serves as a by-product that is used in assisting the generation of $\hat{E_{u,i}}$ in the inference stage.
Figure 3: The proposed MAPLE architecture. MAPLE is trained on Stage 1: Explanation Generation and Stage 2: Aspect Recommendation as a unified architecture. During inference, MAPLE exploits the trained Aspect-Recommendation Model (Rec. Model) in Stage 2 to predict an aspect distribution $\hat{c}_{u,i}$ and obtain an aspect signal $\hat{c}$, and then feed it back to Stage 1 for generation.
@@ 3.2 Multi-Aspect Review Segmentation
Instead of predicting aspect-term importance, as done by ExpansionNet and UARM, our model takes a broader perspective by learning aspect topics and extracting aspect terms from within these topics based on ID information. This approach prevents the model from overfitting to the characteristics of individual users or items, avoiding overly generic explanations that are conditioned on ID alone. By introducing aspect prompts $c$, the model focuses on learning the fine-grained aspect terms of the particular category, while the ID embeddings are learned as selectors for relevant aspects.
We employ an automated sentiment-analysis pipeline to extract sentiment tuples (aspect term, opinion, sentiment, category) from reviews. Concretely, we use a fine-tuned sentiment analysis model (Raffel et al., 2020) to extract sentiment tuples within the reviews and discard the reviews without tuples. We then use a zero-shot classifier (Lewis et al., 2020) to assign each extracted aspect term with an aspect category (for details on model and aspect inventory, see Appendix A.1 and A.3). Finally, for every user-item pair, we check if there are multiple mined aspect terms under the same aspect category; if yes, we choose the tuple with an arbitrary select strategy, e.g., choose the one that is the longest, and add the associated tuple into the training data. Figure 2 illustrates an example of a review and its extracted sentiment tuples, Since we take one tuple for each category, this review yields two tuples. For the "Gluten-Free, Vegan, Vegetarian" category, we take ("vegan pizza", …) since it has a longer aspect term compared to "pie"; for the "Seafood" category, we take ("chowder", …).
3.3 Stage 1: Explanation Generation
@@
We then train MAPLE to condition the user-item-aspect signals for text generation. We minimize the negative log-likelihood loss $L_{T}$, where $\mathcal{T}$ denotes the training set; $E_{u,i,c}$ denotes the explanation segment for the user-item-aspect pair; $c_{3+t}^{e_{t}}$ is offset by 3 to accommodate the user, item, and aspect prompt tokens. As discussed in Section 3.2, rather than learning just one explanation for a user-item pair, a user-item can now pair with different aspect categories $c_{i}$ depending on the categories labeled from the ground-truth review.
@@
3.4 Stage 2: Aspect Recommendation
During training, MAPLE takes a user-item pair and the auxiliary aspect category. However, in the inference stage, the aspect information may be absent. To close the gap between training and inference, we design an auxiliary task, called aspect recommendation, to recover the relationship between IDs and their associated aspect categories. The input of the aspect-recommendation task is a pair of user, item IDs, the output is the predicted aspect category probabilities $\hat{c_{u,i}}\in\mathbb{R}^{n_{\text{aspect}}}$. We implement the model with an MLP architecture:
@@ 3.4 Stage 2: Aspect Recommendation
In equation 2, $\mathbf{W}_{h}\in\mathbb{R}^{h\times 2d}$ and $\mathbf{b}_{h}\in\mathbb{R}^{h}$ are the weights and biases of fully-connected layer(s),   $\mathbf{W}_{c}\in\mathbb{R}^{n_{\text{aspect}}\times h}$ and $\mathbf{b}_{c}\in\mathbb{R}^{n_{\text{aspect}}}$ are the weights and biases of the output classifier layer, and $\delta$ and $\sigma$ are the activation functions ReLU and Sigmoid, respectively.
However, the aspect-category label exhibits a very skewed, long-tail distribution (Figure 5), which poses a challenge in training the aspect-recommendation model, since a conventional classifier easily overfits the head classes. We employ the distribution-balanced loss function (Wu et al., 2020) to address this issue, effectively capturing aspect preferences from ID information rather than label distribution. It is formalized as:
@@
$\Delta_{i}=z^{k}_{i}-\nu_{i}$
In equation 3.4, $\mathcal{L}_{DB}(x^{k},y^{k})$ represents the distribution-balanced loss function for the $k$-th sample; $\hat{r}^{k}_{i}$ represents the sampling weighting factor for the $i$-th aspect class of the $k$-th sample; $y^{k}_{i}$ is the ground truth label for the $i$-th class of the $k$-th sample; $z^{k}_{i}$ is the logit for the $i$-th class of the $k$-th sample; $\nu_{i}$ is the class bias for $i$-th class, and $\lambda$ is a hyperparameter regularizing sigmoid’s gradient issue.
In this stage, we add the recommendation loss weighted by $\alpha$ to the total loss, giving ID embeddings selector ability. Formally,
$L=L_{T}+\alpha L_{DB}$
We conducted preliminary tuning and found that $\lambda=1$ yielded stable training, while larger values (e.g., $\lambda=5$) led to underfitting. Thus, we use $\lambda=1$ in all the reported experiments.
3.5 Inference Stage
Given user and item IDs, we first infer stage 2’s aspect-recommendation model and trim the predicted aspect distribution $\hat{c}_{u,i}$ to leave only the top 5, and then we sample $K(K\leq 5)$ aspects with replacement from the trimmed distribution, using the predicted probabilities as sample weights. The $K$ embeddings are then fused (by taking their mean) to form the aspect signal at position $\hat{c}$, as shown in Figure 3 (b). The trained transformer is thus conditioned on $u,i,\hat{c}$ to generate a well-rounded explanation $\hat{E}_{u,i}$.